OpenStudy (anonymous):

Give an example of a function with both a removable and a non-removable discontinuity.

OpenStudy (anonymous):

@zepdrix I NEEED YOUUUUUUUUUU:))))

zepdrix (zepdrix):

\[\Large y=\frac{x^2-4}{x-2}\]We could simplify this function a bit,\[\Large y=\frac{(x-2)(x+2)}{x-2}\]Which simplifies to,\[\Large y=x+2\]This represents a straight line if we were to graph it. But we still can't use the point x=2 because it causes a zero in the denominator of the original form. So this is an example of a `removable discontinuity`.

zepdrix (zepdrix):

If you're asked to come up with an example, the easiest thing to do is just make a piece-wise function.So for this function, we would graph y=x in all places except x=1. At x=2, the function gives us y=2.